OpenStudy (anonymous):

how to simplify this division radical 3 squared 20 / 2 ?

OpenStudy (dumbcow):

\[\frac{\sqrt{3^{2}}20}{2} = 3*10 = 30\] OR \[\frac{\sqrt{3^{2}*20}}{2}=\frac{\sqrt{3^{2}}\sqrt{4}\sqrt{5}}{2} = 3\sqrt{5}\]